OpenStudy (anonymous):

Last fall, a gardener planted 63 iris bulbs. She found that only 57 of the bulbs bloomed in the spring. Find the empirical probability that an iris bulb of this type will bloom. Give answer as a fraction in lowest terms. How many of the bulbs should she plant next fall if she would like at least 60 to bloom? I see that the first part is 57/63 = 19/21. But I think I'm doing the formula for the 2nd part wrong? I'm doing 19 x (60/21) and getting 54.

OpenStudy (cwrw238):

19/21 * x = 60 x = 60 * 21 ------ = 67 19

Directrix (directrix):

Agree with part one on empirical probability. Part Two 19/21 = 60/x where x is the number to plant to get 60 to bloom 19x = 60 *21 19x = 1260 x = 66.315 -->Plant 67 bulbs